GridPP PMB Meeting 821

GridPP PMB Meeting 821 (6 June 2022)

David Britton, Sam Skipsey (minutes), Andrew McNab, Steve Lloyd, Tony Doyle, Davide Costanzo, Jon Hays, Roger Jones, Katy Ellis, David Colling, David Kelsey, Alastair Dewhurst, Andrew Sansum
Apologies: Pete Clarke, Pete Gronbech, Tony Cass

GridPP 48

RJ in contact with UoCumbria to confirm dates + details. DB has a draft announcement, and SS a registration page waiting on confirmation.
AD to look into sponsorship + list of topics from RAL speakers.

Vision Document (GridPP7)

The vision document from DB was discussed. Potential considerations involving the duration of a future GridPP project, and the interaction with Run 4 timing, and future architectural decisions by the WLCG experiments were discussed.
DB agreed to add suitable wording regarding technologies to the document.

SI-0 (Tier-1 and Technical Meetings)

AD is currently scheduling 3 TMs (Rucio/Dirac, Perfsonar, NewOSes)
AD reported on the current recruitment status at the Tier-1.
Tier-1 had an additional 2hr d/t for ANTARES on Monday morning for final upgrades to networking, successful.
Over the long weekend, operations were stable except for WebDAV tests, due to grossly asymmetric load on one server. James W mitigated the effects for ATLAS, but other VOs were affected. Long term solution is to move from DNS round-robin alias to Xrootd redirection service.
Procurement for next tranche is being discussed with partners. Prices look ~same or slightly more expensive than last year so far.


DCos reports RAL ended the last accounting period @ Pledge level (good!)
Tier-2s – xrootd-related issue with Glasgow on the 2nd (resolved); EDCF resolving an IPv6 issue

SI-2 (CMS)

KE reports RAL WebDAV issue (see SI-0) did cause CMS test failures on multiple days.
Also, ongoing multihop transfer related issue with RAL + FTS, due to bug in FTS (resolved in new release to be installed). Intermediate copies of files sometimes need manually deleted as they cannot be overwritten by new transfers.
Planning for the future test of Tape performance until post FTS fix.
KE also has some feedback on Shoveler tests from CERN endpoint (things look good on their end); waiting for CERN monitoring to be up before rolling out more endpoints.

SI-3 (LHCb)

AMcN reports some clean up of files happening with LHCb.
Several RAL tickets are now closed(good!), down to just 3.
Tier-2s fine.

SI-4 (Ops Report)

SS reports Ops cancelled last week due to long weekend; other meetings with Tier-2s that did happen have NTR.

SI-5 (MB)


SI-6 (External)

SEAG meeting on Wednesday, DB and other attending.

Review of Actions

782.4 DCos – Investigate VAC migration plan for Birmingham. [Ongoing – raised at UK Cloud Support meeting. No formal request as yet.] – follow up

800.5 AD – Arrange in person DIRAC/Rucio meeting at IC (Jan22). [on-going] – AD plans to re-establish virtual meetings and then set up an in person meeting. AD and DC to try and arrange a date.

818.4 – JS and SS to produce and add the PMB minutes to the CERN website every Monday morning when the Agenda is circulated to the PMB (Ongoing)

818.5 – DB/SS Group to formalise VOs to be added to the approved list.